# InfiniteHBD: Building Datacenter-Scale High-Bandwidth Domain for LLM with Optical Circuit Switching Transceivers

### v3

@@ 6.2. HBD Fault Resilience
GPU waste. Apart from faulty GPUs, issues such as fragmentation, topology disconnections, and bandwidth degradation can render healthy GPUs wasted. The GPU waste ratio quantifies the number of wasted GPUs under different fault scenarios. Figure 9 illustrates GPU waste ratios over production trace, while Figure 10 depicts the GPU waste ratio as node fault ratio vary.
@@
Table 5. Optimal parallelism strategies for maximize MFU of GPT-MoE under varying GPU numbers.
In these scenarios, InfiniteHBD ($K=3$) achieves near-zero GPU waste ratio, and outperforming all other architectures. Especially, the waste ratio for InfiniteHBD ($K=2$) remains almost identical to InfiniteHBD ($K=3$), allowing one bundle of OCSTrx to be saved for clusters with low fault rates. NVL-36 and NVL-72 typically experience an 11% waste ratio for TP sizes of 16 or larger, as $1/9$ of GPUs are reserved for redundant backups. NVL-576 has less fragmentation, benefiting from its larger size. TPUv4 performs well at low fault ratios and small TP sizes, but significantly degrades with larger TP sizes due to its coarse $4^{3}$ cube-based resource management, which amplifies the fault explosion radius. To sum up, InfiniteHBD demonstrates the strongest fault resilience among all architectures.
Maximum job supported. In fixed-size clusters, large job must pause when the available GPUs drop below the required count. Faced with same fault rate, cluster with lower GPU waste ratio can support larger job scales. Figure 11 shows the maximum job scale supported for various HBD architectures cluster with 9,792-GPU, simulated with the fault traces normalized for 4-GPU nodes. InfiniteHBD ($K=2$ or $K=3$) and NVL-576 lead in performance, and SiP-Ring exhibits declining efficiency as TP size increases.
Figure 11. Maximal job scale supported by 9,792 GPUs.
Job fault-waiting time. Large job must wait for the repairing when GPU availability falls below the required threshold. This simulations assume the average recovery time in the fault trace as a fixed repair duration. The total wasted time during 348 days is evaluated (Figure 12). For smaller TP sizes (TP-8/TP-16), NVL-36/NVL-72 exhibit the weakest resilience due to their 11% backup overhead. For larger TP sizes (TP-32/TP-64), SiP-Ring and TPUv4 perform worst.
@@ 6.3. Training Performance
This section analyzes the training performance of two representative large models, LLama 3.1-405B (Llama Team, AI @ Meta, 2024) and GPT-MoE (configuration detailed in Appendix §B), under various GPU resource configurations and parallelism strategies. The simulation results validate the practical applicability of the InfiniteHBD architecture. In simulations, we model practical TP and EP behaviors: For TP, increasing parallelism splits GEMMs into smaller, less efficient tasks, reducing hardware efficiency (NVIDIA, 2023); for EP, we practically set expert imbalance coefficient at 20%.
@@ 6.4. Communication Efficiency
This section examines the impact of orchestration algorithms on DCN communication efficiency. Experiments were performed on a Fat-Tree architecture, like the setup in  (Gangidi et al., 2024b). As shown in Figure 13(a), the algorithm is not sensitive to cluster size. Therefore, the evaluation is based on TP-32 operations on InfiniteHBD with 8192 GPUs.
  • Baseline: A greedy algorithms, which randomly select nodes from the cluster and use the first permutation that meets the requirements.
@@
Figure 13(b) illustrates the impact of job-scale ratios (job size/total cluster GPUs) on cross-ToR traffic, where node fault ratio is 5%. Baseline consistently results in approximately 10% cross-ToR traffic. In contrast, the Optimized algorithm significantly outperforms the Baseline, reducing cross-ToR traffic to just 1.72% even at a 90% job-scale ratio. Figure 13(c) explores the sensitivity to node faults, with the job scale ratio fixed at 85%. The Baseline shows a linear increase of cross-ToR traffic, while the Optimized algorithm sustains near-zero cross-ToR traffic for fault ratios under 7%.
@@ 6.5. Cost and Power Analysis
The breakdown analysis of each architecture is provided in the Appendix §F. Based on this, the cost and power consumption are normalized according to GPU count and per-GPU bandwidth. As depicted in Table 6, InfiniteHBD exhibits the lowest interconnect cost per GPU per GBps. Under the $K=2$ configuration, its cost is only 62.84% of Google TPUv4 and 30.86% of the NVIDIA GB200 NVL-36/72, with minimal power consumption. This efficiency is primarily attributed to the avoidance of centralized switches. TPUv4 ranks second in interconnect cost and lowest in power consumption, achieved by reducing optical module use and per-port OCS costs. The NVL series has higher interconnect costs and power consumption due to its fully-connected topology and high-cost NVLink Switches. Notably, NVL-576 incurs the highest cost and power consumption due to its multilayer nonconvergent topology, which increases optical module expenses and requires more NVLink Switches.
@@
Simulations on a 10K-GPU cluster using the TP-32 configuration evaluate GPU availability under varying fault ratios across different architectures. The variation in aggregate cost for different HBD architectures under varying node fault ratios is illustrated in Figure 13(d). InfiniteHBD consistently exhibits the lowest aggregate cost. Furthermore, when the fault ratio is below 12.1%, the aggregate cost of InfiniteHBD ($K=2$) is less than that of InfiniteHBD ($K=3$), suggesting that ($K=2$) is the optimal design for most scenarios.
@@ 7. Discussion
AllToAll communication. Ring topology in InfiniteHBD struggles with AllToAll communication (e.g., EP), exhibiting poor performance at $O(p^{2})$, where $p$ is the group size. This can be improved by linking backup lines to nodes indexed at $n\pm 2^{i}$ instead of $n\pm i$ and Applying the Binary Exchange algorithm, and reduce time complexity to $O(p\log_{2}p)$. During the algorithm, OCSTrx need to connect to different GPUs with runtime switching, since OCSTrx switches in under $1ms$, reconfiguration can be overlapped with computation. For $K=2$ InfiniteHBD designs, performance matches the ideal Bruck algorithm (Bruck et al., 1997) when $p<8$. However, this design introduces complexities in construction, failover, and orchestration, and necessitates GPU routing capabilities. So it is not applied.
Simulation Scale. Simulations using real 10,000 GPU fault traces, as detailed in §6.2, were conducted on a cluster comprising 9,792 GPUs. The reason is that the simulation’s GPU count must be less than the total GPUs in the fault trace, and 9,792 is the largest number divisible by 576 and less than 10,000. This configuration allows the entire cluster to be divided into 17 NVL-576 units for the simulation.
Multi-dimension parallelism. InfiniteHBD is optimized for single-dimension parallelism. To support multi dimensional communication, two approaches are viable. 1) Independent Interconnects: Each OCSTrx bundle includes multiple OCSTrx units (e.g., 4 or 8), then link each of the unit to a separate inter-host topology. This isolates parallel dimensions but results in fixed bandwidth per dimension, leading to inefficiencies. 2) Time-Division multi-dimension: Main and backup lines of OCSTrx can be used to form separated inter-host topology, rapidly switching between them can support multi-dimension parallelism. However, this introduces complexity in managing multi-dimensional overlap and reduces the fault tolerance of InfiniteHBD.
Single-Job vs. Multi-Job. Existing studies explore multi-job scheduling in GPU clusters (Weng et al., 2022; Li et al., 2023b). Deploying certain small jobs, such as inference tasks, can mitigate GPU fragmentation. However, given the shortage of GPU in LLM training, any idle GPU—whether repurposed for small job or not—is undesirable. Thus, InfiniteHBD prioritizes single-job execution for simplicity.
OCS vs. EPS. OCSTrx enables multi-path selection, a feature also achievable with Electronic Packet Switching (EPS). For example, inter-host topology of InfiniteHBD can be implemented using UBB 2.0-based servers by adding external optical interfaces to switches in server. However, this would require twice optical modules and numerous high-throughput switching chips for the entire system, significantly increasing cost and power consumption compared to OCSTrx.
@@ 8. Related Work
HBD Architectures. HBDs are crucial for enabling communication intensive parallelism strategies (TP/EP) for LLM training. NVIDIA DGX SuperPOD (NVIDIA, 2024a) and GB200 NVL series (NVIDIA, 2024b) use any-to-any electrical switching, delivering high performance but suffering from high costs, scalability limitations, and fragmentation. In contrast, direct interconnect HBDs like Dojo (Tesla, 2024), TPUv3 (Jouppi et al., 2020), and SiP-Ring (Khani et al., 2021) improve scalability but have a large fault explosion radius. TPUv4 (Jouppi et al., 2023) and TPUv5p (Google, 2024) attempts a middle ground but still lacks full node-level fault isolation. InfiniteHBD introduces a novel architecture that reduces cost, improves scalability, minimizes fragmentation, enhances fault isolation, and dynamically supports TP.
AI DCN Architectures. MegaScale (Jiang et al., 2024a) and Meta’s (Gangidi et al., 2024b) AI DC use Clos-based topologies, while Rail-Optimized (NVIDIA, 2021) and Rail-Only (Wang et al., 2024) architectures optimize for LLM traffic patterns. Alibaba HPN (Qian et al., 2024) enhances fault tolerance with a dual-plane design. InfiniteHBD is compatible with all of them on LLM-training.
OCS Technologies. OCS enables dynamic topology reconfiguration in datacenters (Urata et al., 2022b; Jouppi et al., 2023; Liao et al., 2025). MEMS OCS-based switch supports high port counts (Urata et al., 2022a; Calient.AI, 2024), while silicon photonics (SiPh) achieves lower latency and cost (Hauffe and Petermann, 2006). This work proposes a SiPh-based OCS transceiver (OCSTrx), constructing an interconnect fabric without centralized switches.
@@
AI Job Schedulers. Schedulers such as  (Xiao et al., 2018; Mahajan et al., 2020; Gu et al., 2019; Jiang et al., 2020; Peng et al., 2019; Qiao et al., 2021) aim to improve GPU utilization. However, they exhibit dual limitations: their designs are premised on non-reconfigurable network, while also failing to consider job scheduling within HBD for optimizing traffic patterns in DCN. This work proposes a HBD-DCN orchestration algorithm based on reconfigurable networks to address these limitations.
@@ 9. Conclusion
In this paper, we propose InfiniteHBD, a novel HBD design that supports datacenter scale, dynamic TP group size and near-ideal fault explosion radius. InfiniteHBD is built upon a novel design of optical transceivers interaged with SiPh-based OCSTrx, a reconfigurable K-Hop Ring topology and a HBD-DCN orchestration algorithm to leverage the capabilities of the new hardware. Using real fault trace of 10K GPU cluster and the in-house simulator, we demonstrate that InfiniteHBD achieves GPU utilization close to the ideal model during faults, delivers superior cost and energy efficiency compared to existing designs, and provides effective control over cross-ToR DCN traffic. We believe InfiniteHBD provides an efficient scaling solution for HBD, which offers new insights for the infrastructure of the next generation Trillion-Parameters LLM training.
@@ Appendix A Production Fault Trace
The production fault trace was collected from an 8-GPU node pretrain cluster with 10K GPUs over a period of 348 days. The trace includes details such as fault start time, fault end time, and the ID of the faulty node. Figure 14(a) and Figure 14(b) provide a macro-level overview of the production fault trace. On average, the ratio of faulty 8-GPU nodes at any given time is $2.66\%$, with a p99 value of $7.92\%$.
@@
Since most of failure events are GPU faults, we normalized the trace of 8-GPU nodes to generate 4-GPU nodes trace. Assuming that the fault rates of GPUs are i.i.d. with a fault probability of $p$ for each GPU, and considering that a node is deemed faulty if any GPU within it fails, the fault rate of an 8-GPU node is calculated as:
$P_{fault}(8\text{-GPU})=1-(1-p)^{8}=2.66\%.$
From this, we derive $p=0.34\%$. The fault rate for a 4-GPU node is then:
$P_{fault}(4\text{-GPU})=1-(1-p)^{4}=1.34\%.$
The fault event of 4-GPU node is generate with Bayesian Equation, as:
@@
$=\frac{1\times 1.34\%}{2.66\%}=50.38\%$
Thus, whenever a fault occurs in an 8-GPU node in the original trace, each of the two corresponding 4-GPU nodes at the same location has a $50.38\%$ probability of fault. This method is used to convert the traces.
As node faults are i.i.d., the simulator linearly maps the fault trace to different network architectures.

### v4

@@ 6.2. HBD Fault Resilience
GPU waste. Apart from faulty GPUs, issues such as fragmentation, topology disconnections, and bandwidth degradation can render healthy GPUs wasted. The GPU waste ratio quantifies the number of wasted GPUs under different fault scenarios. Figure 13 illustrates GPU waste ratios over production trace, while Figure 14 depicts the GPU waste ratio as node fault ratio varies.
@@
Table 5. Optimal parallelism strategies for maximizing MFU of GPT-MoE under varying GPU numbers.
In these scenarios, InfiniteHBD ($K=3$) achieves near-zero GPU waste ratio, and outperforms all other architectures. Especially, the waste ratio for InfiniteHBD ($K=2$) remains almost identical to that of InfiniteHBD ($K=3$), allowing one bundle of OCSTrx to be saved for clusters with low fault rates. NVL-36 and NVL-72 typically experience an 11% waste ratio for TP sizes of 16 or larger, as $1/9$ of GPUs are reserved for redundant backups. NVL-576 has less fragmentation, benefiting from its larger size. TPUv4 performs well at low fault ratios and small TP sizes, but significantly degrades with larger TP sizes due to its coarse $4^{3}$ cube-based resource management, which amplifies the fault explosion radius. To sum up, InfiniteHBD demonstrates the strongest fault resilience among all architectures.
Maximum job supported. In fixed-size clusters, large jobs must pause when the available GPUs drop below the required count. Faced with the same fault rate, clusters with lower GPU waste ratio can support larger job scales. Figure 15 shows the maximum job scale supported for various HBD architectures cluster with 2,880-GPU, simulated with the fault traces normalized for 4-GPU nodes. InfiniteHBD ($K=2$ or $K=3$) and NVL-576 lead in performance, and SiP-Ring exhibits declining efficiency as TP size increases.
Figure 15. Maximal job scale supported by 2,880 GPUs.
Job fault-waiting time. Large jobs must wait for repairs when GPU availability falls below the required threshold. These simulations assume the average recovery time in the fault trace as a fixed repair duration. The total wasted time during 348 days is evaluated (Figure 16). For smaller TP sizes (TP-8/TP-16), NVL-36/NVL-72 exhibit the weakest resilience due to their 11% backup overhead. For larger TP sizes (TP-32/TP-64), SiP-Ring and TPUv4 perform worst.
@@ 6.3. Training Performance
This section analyzes the training performance of two representative large models, LLama 3.1-405B (Grattafiori et al., 2024) and GPT-MoE (configuration detailed in Appendix §B), under various GPU resource configurations and parallelism strategies. The simulation results validate the practical applicability of the InfiniteHBD architecture. In simulations, we model practical TP and EP behaviors: For TP, increasing parallelism splits GEMMs into smaller, less efficient tasks, reducing hardware efficiency (NVIDIA, 2023); for EP, we practically set expert imbalance coefficient at 20%.
@@ 6.4. Communication Efficiency
This section examines the impact of orchestration algorithms on DCN communication efficiency. Experiments were performed on a Fat-Tree architecture, like the setup in  (Gangidi et al., 2024). As shown in Figure 17(a), the algorithm is not sensitive to cluster size. Therefore, the evaluation is based on TP-32 operations on InfiniteHBD with 8192 GPUs.
  • Baseline: A greedy algorithm, which randomly selects nodes from the cluster and uses the first permutation that meets the requirements.
@@
Figure 17(b) illustrates the impact of job-scale ratios (job size/total cluster GPUs) on cross-ToR traffic, where node fault ratio is 5%. The Baseline consistently results in approximately 10% cross-ToR traffic. In contrast, the Optimized algorithm significantly outperforms the Baseline, reducing cross-ToR traffic to just 1.72% even at a 90% job-scale ratio. Figure 17(c) explores the sensitivity to node faults, with the job scale ratio fixed at 85%. The Baseline shows a linear increase of cross-ToR traffic, while the Optimized algorithm sustains near-zero cross-ToR traffic for fault ratios under 7%.
@@ 6.5. Cost and Power Analysis
The breakdown analysis of each architecture is provided in the Appendix §F. Based on this, the cost and power consumption are normalized according to GPU count and per-GPU bandwidth. As depicted in Table 6, InfiniteHBD exhibits the lowest interconnect cost per GPU per GBps. Under the $K=2$ configuration, its cost is only 62.84% of Google TPUv4 and 30.86% of the NVIDIA GB200 NVL-36/72, with minimal power consumption. This efficiency is primarily attributed to the avoidance of centralized switches. TPUv4 ranks second in interconnect cost and lowest in power consumption, achieved by reducing optical module use and per-port OCS costs. The NVL series has higher interconnect costs and power consumption due to its fully-connected topology and high-cost NVLink Switches. Notably, NVL-576 incurs the highest cost and power consumption due to its multilayer non-blocking topology, which increases optical module expenses and requires more NVLink Switches.
@@
Simulations on a 3K-GPU cluster using the TP-32 configuration evaluate GPU availability under varying fault ratios across different architectures. The variation in aggregate cost for different HBD architectures under varying node fault ratios is illustrated in Figure 17(d). InfiniteHBD consistently exhibits the lowest aggregate cost. Furthermore, when the fault ratio is below 12.1%, the aggregate cost of InfiniteHBD ($K=2$) is less than that of InfiniteHBD ($K=3$), suggesting that ($K=2$) is the optimal design for most scenarios.
@@ 7. Discussion
AllToAll communication. Ring topology in InfiniteHBD struggles with AllToAll communication (e.g., EP), exhibiting poor performance at $O(p^{2})$, where $p$ is the group size. This can be improved by linking backup lines to nodes indexed at $n\pm 2^{i}$ instead of $n\pm i$ and applying the Binary Exchange algorithm, reducing time complexity to $O(p\log_{2}p)$. During the algorithm, OCSTrx needs to connect to different GPUs with runtime switching; since OCSTrx switches in 60-80 $\mu s$, reconfiguration can be overlapped with computation. For $K=2$ InfiniteHBD designs, performance matches the ideal Bruck algorithm (Bruck et al., 1997) when $p<8$. However, this design introduces complexities in construction, failover, and orchestration, and necessitates GPU routing capabilities. Therefore, it is not applied. We provide a detailed theoretical discussion on how InfiniteHBD supports AllToAll communication in Appendix §G.
Simulation Scale. Simulations using real fault traces from 3,200 GPUs, as detailed in §6.2, were conducted on a cluster comprising 2,880 GPUs. This is because the simulation’s GPU count must be less than the total GPUs in the fault trace, and 2,880 is the largest number divisible by 576 and less than 3,200. This configuration allows the entire cluster to be divided into 5 NVL-576 units for the simulation.
Multi-dimension parallelism. InfiniteHBD is optimized for single-dimension parallelism. To support multi-dimensional communication, two approaches are viable. 1) Independent Interconnects: Each OCSTrx bundle includes multiple OCSTrx units (e.g., 4 or 8), then links each of the units to a separate inter-host topology. This isolates parallel dimensions but results in fixed bandwidth per dimension, leading to inefficiencies. 2) Time-Division multi-dimension: Main and backup lines of OCSTrx can be used to form separated inter-host topologies, rapidly switching between them can support multi-dimensional parallelism. However, this introduces complexity in managing multi-dimensional overlap and reduces the fault tolerance of InfiniteHBD.
Single-Job vs. Multi-Job. Existing studies explore multi-job scheduling in GPU clusters (Weng et al., 2022; Li et al., 2023b). Deploying certain small jobs, such as inference tasks, can mitigate GPU fragmentation. However, given the shortage of GPUs in LLM training, any idle GPU—whether repurposed for small jobs or not—is undesirable. Thus, InfiniteHBD prioritizes single-job execution for simplicity.
OCS vs. EPS. OCSTrx enables multi-path selection, a feature also achievable with Electronic Packet Switching (EPS). For example, the inter-host topology of InfiniteHBD can be implemented using UBB 2.0-based servers by adding external optical interfaces to switches in servers. However, this would require twice the number of optical modules and numerous high-throughput switching chips for the entire system, significantly increasing cost and power consumption compared to OCSTrx.
@@ 8. Related Work
HBD Architectures. HBDs are crucial for enabling communication-intensive parallelism strategies (TP/EP) for LLM training. NVIDIA DGX SuperPOD (NVIDIA, 2024a) and GB200 NVL series (NVIDIA, 2024b) use any-to-any electrical switching, delivering high performance but suffering from high costs, scalability limitations, and fragmentation. In contrast, direct interconnect HBDs like Dojo (Tesla, 2024), TPUv3 (Jouppi et al., 2020), and SiP-Ring (Khani et al., 2021) improve scalability but have a large fault explosion radius. TPUv4 (Jouppi et al., 2023) and TPUv5p (Google, 2024) attempt a middle ground but still lacks full node-level fault isolation. InfiniteHBD introduces a novel architecture that reduces cost, improves scalability, minimizes fragmentation, enhances fault isolation, and dynamically supports TP.
AI DCN Architectures. MegaScale (Jiang et al., 2024a) and Meta’s (Gangidi et al., 2024) AI DC use Clos-based topologies, while Rail-Optimized (NVIDIA, 2021) and Rail-Only (Wang et al., 2024) architectures optimize for LLM traffic patterns. Alibaba HPN (Qian et al., 2024) enhances fault tolerance with a dual-plane design. InfiniteHBD is compatible with all of them for LLM training.
OCS Technologies. OCS enables dynamic topology reconfiguration in datacenters (Urata et al., 2022; Jouppi et al., 2023; Liao et al., 2025). A MEMS OCS-based switch supports high port counts (Urata et al., 2022; Calient.AI, 2024), while silicon photonics (SiPh) achieves lower reconfiguration latency and cost (Hauffe and Petermann, 2006). This work proposes a SiPh-based OCS transceiver (OCSTrx), which constructs an interconnect fabric without centralized switches.
@@
AI Job Schedulers. Schedulers such as  (Xiao et al., 2018; Mahajan et al., 2020; Gu et al., 2019; Jiang et al., 2020; Peng et al., 2019; Qiao et al., 2021) aim to improve GPU utilization. However, they exhibit dual limitations: their designs are premised on a non-reconfigurable network, while also failing to consider job scheduling within HBD for optimizing traffic patterns in DCN. This work proposes a HBD-DCN orchestration algorithm based on reconfigurable networks to address these limitations.
@@ 9. Conclusion
In this paper, we propose InfiniteHBD, a novel HBD design that supports datacenter scale, dynamic TP group size and near-ideal fault explosion radius. InfiniteHBD is built upon a novel design of optical transceivers integrated with SiPh-based OCSTrx, a reconfigurable K-Hop Ring topology and a HBD-DCN orchestration algorithm to leverage the capabilities of the new hardware. Using a real fault trace of 3K-GPU cluster and the in-house simulator, we demonstrate that InfiniteHBD achieves GPU utilization close to the ideal model during faults, delivers superior cost and energy efficiency compared to existing designs, and provides effective control over cross-ToR DCN traffic. We believe InfiniteHBD provides an efficient scaling solution for HBD, which offers new insights for the next-generation infrastructure for training trillion-parameter LLMs.
Acknowledgements.
We thank our shepherd Yiting Xia and anonymous reviewers for their constructive feedback. We are grateful to Juston Zhu, Zhijie Chen, Ziyao Shen, Jie Chen, and Edmund Xu from Lightelligence Pte. Ltd. for their assistance in hardware evaluation. We also thank Instruments and Electronics (Shanghai) Associates (INESA) for providing a GPU cluster optical interconnect testing platform. This work is supported by Beijing Municipal Science and Technology Project No. Z241100004224023, National Natural Science Fund for the Excellent Young Scientists Fund Program (Overseas), and Peking University startup fund.
APPENDICES
Appendices are supporting material that has not been peer-reviewed.
@@ Appendix A Production Fault Trace
The production fault trace was collected from a 3K-GPU cluster dedicated to pretrain with 8-GPU nodes during a period of 348 days. The trace includes details such as fault start time, fault end time, and the ID of the faulty node. Figure 18(a) and Figure 18(b) provide a macro-level overview of the production fault trace. On average, the ratio of faulty 8-GPU nodes at any given time is $2.33\%$, with a P99 value of $7.22\%$.
@@
Since most of the failure events are GPU faults, we normalized the trace of 8-GPU nodes to generate 4-GPU nodes trace. Assuming that the fault rates of GPUs are i.i.d. with a fault probability of $p$ for each GPU, and considering that a node is deemed faulty if any GPU within it fails, the fault rate of an 8-GPU node is calculated as:
$P_{fault}(8\text{-GPU})=1-(1-p)^{8}=2.33\%.$
From this, we derive $p=0.29\%$. The fault rate for a 4-GPU node is then:
$P_{fault}(4\text{-GPU})=1-(1-p)^{4}=1.17\%.$
The fault event trace for 4-GPU nodes is generated using Bayesian Equation as follows:
@@
$=\frac{1\times 1.17\%}{2.33\%}=50.21\%$
Thus, whenever a fault occurs in an 8-GPU node in the original trace, each of the two corresponding 4-GPU nodes at the same location has a $50.21\%$ probability of failure. This method is used to convert the traces.
As node faults are assumed to be i.i.d., the simulator linearly maps the fault trace onto different network architectures.